Please help me solve. I am confused.
The table shows how much money different people made at various jobs. Fill in the blanks.
#of hours Total $
0 4
1 8
2 12

Unit rate of change: _______________________
Linear? _____yes
Dependent Variable: y=__Total dollar amt per hour
Independent Variable: x _#of hours
Starting Point: _4___ __________________
Equation: ________________________________
Check

Answer by Theo(13342)

when they work 0 hours, they make 4 dollars.
when they work 1 hour, they make 8 dollars.
when they work 2 hours, they make 12 dollars.

looks like they make 4 dollars an hour at the beginning of the hour.

the equation would be y = 4 * x + 4

when x = 0, y = 0 + 4 = 4
when x = 1, y = 4 + 4 = 8
when x = 2, y = 8 + 4 = 12

answers to your questions:

Unit rate of change: __________4 collars an hour_____________
Linear? _____yes
Dependent Variable: y=__total dollars made !!!!!
Independent Variable: x _#of hours
Starting Point: _4___ __________________
Equation: ____y = 4x + 4____________________________
Check

maybe they make 4 dollars just to show up and then make 4 dollars an hour on top of that.

that might make more sense.

the objective was to get you to write an equation.
not sure the problem was posed very well, because it is confusing.
stil, the answers and the equation are what they are looking for.

your answer to the dependent variable was that y represents the total amount of money spent.

if they work 0 hours, they make 4 dollars.
if they work 1 hour, they make 8 dollars.
if they work 2 hours, they make 12 dollars.,
etc.